Problem 2: For a mixture of benzene and toluene (XA-0.33) at l 00 °C, determine the system P and assumptions made in solving this problem. List Problem 3:
Answer #1

Antoine equation is given by

log10(Psat[mmHg])=A-B/(T[oC]+C)

A B C
toluene 6.95087 1342.31 219.187
benzene 6.87987 1196.76 219.161

At T = 100 C,

saturation pressure of toluene,

P_{T}^{sat}= 10^{6.95087-\frac{1342.31}{100+219.187}} = 556.5\ mm\ Hg

For benzene,

P_{B}^{sat}= 10^{6.87987-\frac{1196.76}{100+219.161}} = 1349.47\ mm\ Hg

Assuming that the mixture is Ideal and hence follows Raoult's law.

By Raoult's law,

yBP = xBPBsat

yTP = xTPTsat

On adding these equation, we get

yBP + yTP = xBPBsat + xTPTsat

.: P =  xBPBsat + xTPTsat ........................................as yB + yT = 1

.: P = 0.33*1349.47 + (1-0.33)*556.5

.: P = 818.18 mm Hg.......................................................Answer

Now,

yBP = xBPBsat

.: yB = (0.33*1349.47) / (818.18)

.: yB = 0.544287.......................................................Answer